OpenStudy (anonymous):

What is the solution the proportion? 9 4 - = - x 12 A. 13 B. 27 C. 3 D. 17

OpenStudy (anonymous):

WILL MEDAL!

OpenStudy (yanasidlinskiy):

Never "cross multiply". It is utterly silly and useless. Simply solve for the variable. 9/x = 4/12 Simplfy where possible 9/x = 1/3 x cannot be zero, so multiply by that. 9 = x/3 Multiply by 3. x = 27 Done!

OpenStudy (anonymous):

Thank you so much!:D

OpenStudy (yanasidlinskiy):

No problem!:) Anytime!:)